32. You purchased a 12-year Treasury bond with a 7 percent coupon rate. The bond’s asked...

32.

You purchased a 12-year Treasury bond with a 7 percent coupon rate. The bond’s asked yield is currently 5.5 percent. The settlement of the purchase occurred 20 days after the last coupon payment and there are 160 days before the next. If the bond has a par value of $1,000 and interest is paid semiannually. What is the full price of the bond that you pay to the seller?

$1,090.21

$1,130.51

$1,144.12

$1,121.30

$1,134.40
Answer #1

First we need to the find the quoted price, for that we need to put the following values in the financial calculator:

INPUT 12*2=24 5.5/2=2.75 (7%/2)*1,000=35 1,000
TVM N I/Y PV PMT FV
OUTPUT -1,130.51

So, the quoted price = $1,130.51

But when the bond is sold in between the coupon dates, then the actual price paid includes the accrued interest alongwith the quoted price.

Accrued Interest = Coupon payment x [Days since last coupon payment / total days in the period]

= $35 x [20/180] = $3.89

So, Full Price = Clean Price + Accrued Interest = $1,130.51 + $3.89 = $1,134.40

Hence, Option "E" is correct.